# Collaborative Deterministic-Diffusion Model for Probabilistic Urban Spatiotemporal Prediction

### v1

@@ 2.2. Deterministic Prediction
Deterministic prediction of spatiotemporal data is a point estimation approach that assumes a deterministic prediction can be made given the same historical observations. These models are typically trained using loss functions such as MSE or MAE, ensuring that the model learns to predict the conditional mean $\mathbb{E}[y|x]$ to capture the primary patterns. Existing deep learning-based deterministic forecasting models include MLP-based (Shao et al., 2022; Qin et al., 2023; Zhang et al., 2023), CNN-based (Li et al., 2017; Liu et al., 2018; Zhang et al., 2017), and RNN-based (Bai et al., 2019b; Lin et al., 2020; Wang et al., 2017, 2018) architectures, which are commonly used for their computational efficiency. Additionally, GNN-based models (Bai et al., 2019a, a, 2020; Geng et al., 2019; Jin et al., 2023a) have been developed to capture spatial dependencies in graph-structured data, while Transformer-based models (Chen et al., 2022, 2021a; Jiang et al., 2023; Yu et al., 2020) excel at capturing complex dynamic patterns.
@@ 3.3. Evaluation of Probabilistic Prediction
$r_{m}=\frac{1}{N}\sum_{n=1}^{N}\mathbb{1}_{y_{n}\geq\hat{y}_{n}^{\text{low}_{m}}}\cdot\mathbb{1}_{y_{n}\leq\hat{y}_{n}^{\text{high}_{m}}},$
@@
(5) $IS:=\frac{1}{N}\sum_{n=1}^{N}\left[(u_{n}^{\alpha}-l_{n}^{\alpha})+\frac{2}{\alpha}(l_{n}^{\alpha}-y_{n})\mathbb{1}_{y_{n}<l_{n}^{\alpha}}+\frac{2}{\alpha}(y_{n}-u_{n}^{\alpha})\mathbb{1}_{y_{n}>u_{n}^{\alpha}}\right],$
@@ 4.3. Residual Learning via Diffusion Model
(8) $\mathbf{r}_{ta}=\mathbf{x}^{ta}-\mathbb{E}_{\theta}[\mathbf{x}^{ta}|\mathbf{x}^{co}].$
Consequently, the target data $\mathbf{x}^{ta}$ for diffusion models in Eqs. (8), (2), and (3) is replaced by $\mathbf{r}_{ta}$. The residual distribution of urban spatiotemporal data is not independently and identically distributed (i.i.d.) nor does it follow a fixed distribution, such as $\mathcal{N}(0,\mathbf{\sigma})$. Instead, it often exhibits complex spatiotemporal dependence and heterogeneity. So we consider both temporal residual learning and spatial residual learning.
@@
Accordingly, the Eq 13 in the forward process is rewritten as:
@@
And in the denoising process, we sample $\mathbf{r}_{N}^{ta}$ from $\mathcal{N}({\mathbf{Q}},I)$, and denoise it use Eq (2), the computation of $\mu_{\theta}(\mathbf{r}_{n}^{ta},n|\mathbf{x}_{0}^{co})$ in Eq 2 is modified as:
@@ 5. Performance Evaluations
    DyDiffusion can only be applied to grid-format data.
@@ 5.1. Experimental Settings
To evaluate the performance, we employed two deterministic metrics, MAE and RMSE, along with three probabilistic metrics: CRPS, QICE, IS. For QICE, we set the number of QIs, denoted as $M_{\text{QIs}}$, to 10. For IS, we choose a confidence level of 90%, that is, $\alpha=0.1$.
@@ 5.3. Long-Term Prediction
Results of Deterministic Error Metrics.
@@ 5.5. Model Efficiency
We evaluated the training time of our model and all baseline models on the CellularSH dataset, as well as the time required to perform 50 sampling iterations on the test set. Notably, the pretraining time of the mean prediction module in our model is also included in the evaluation. The experimental results are summarized in Table 5. As shown in Table 5, our model demonstrates a significant advantage over existing probabilistic models in terms of training and inference efficiency. This efficiency is particularly valuable for real-world deployment scenarios. It is worth noting that while the CSDI model benefits from the Transformer architecture’s capability to capture spatiotemporal dependencies, its training and sampling time is considerably higher than that of other models, posing challenges for time-sensitive tasks.
@@ 5.6. Analysis of Distribution Alignment
For regions with complex distributions in the CellularSH dataset, we validate the model’s ability to capture multi-modal characteristics by comparing the predicted and true distributions using Kernel Density Estimation (KDE). As shown in Figure 4, the real distribution exhibits significant spatiotemporal multi-modality: Figure 4(a) shows three peaks, which likely represent values at different time points and values under different states at the same time points. CoST successfully captures the three peaks, while CSDI only fits two, demonstrating superior capability in modeling multimodal distributions. In Figure 4(b), the difference between the two peaks reflects strong temporal dynamics in this region. Although both models predict two peaks, the peak spacing of CoST is more consistent with the true distribution. These advantages stem from the model’s design: the diffusion model focuses on residuals, effectively capturing multimodal distribution features, while the deterministic model captures long-term trends and enhances the model’s capacity to model primary dynamic patterns.

### v3

# Collaborative Deterministic-Diffusion Model for Probabilistic Spatiotemporal Prediction

@@ 2.2. Deterministic Prediction
Deterministic prediction of spatiotemporal data is a point estimation approach that assumes a deterministic prediction can be made given the same historical observations. These models are typically trained using loss functions such as MSE or MAE, ensuring that the model learns to predict the conditional mean $\mathbb{E}[y|x]$ to capture the primary patterns. Existing deep learning-based deterministic forecasting models include MLP-based (Shao et al., 2022; Qin et al., 2023; Zhang et al., 2023), CNN-based (Li et al., 2017; Liu et al., 2018; Zhang et al., 2017), and RNN-based (Bai et al., 2019b; Lin et al., 2020; Wang et al., 2017, 2018) architectures, which are commonly used for their computational efficiency. Additionally, GNN-based models (Bai et al., 2019a, 2020; Geng et al., 2019; Jin et al., 2023a) have been developed to capture spatial dependencies in graph-structured data, while Transformer-based models (Chen et al., 2022, 2021a; Jiang et al., 2023; Yu et al., 2020) excel at capturing complex dynamic patterns.
@@ 3.3. Evaluation of Probabilistic Prediction
$r_{m}=\frac{1}{N}\sum_{n=1}^{N}\mathds{1}_{y_{n}\geq\hat{y}_{n}^{\text{low}_{m}}}\cdot\mathds{1}_{y_{n}\leq\hat{y}_{n}^{\text{high}_{m}}},$
@@
(5) $\text{IS}:=\frac{1}{N}\sum_{n=1}^{N}\left[(u_{n}^{\alpha}-l_{n}^{\alpha})+\frac{2}{\alpha}(l_{n}^{\alpha}-y_{n})\mathds{1}_{y_{n}<l_{n}^{\alpha}}+\frac{2}{\alpha}(y_{n}-u_{n}^{\alpha})\mathds{1}_{y_{n}>u_{n}^{\alpha}}\right],$
@@ 4.3. Residual Learning via Diffusion Model
(8) $\mathbf{r}^{ta}=\mathbf{x}^{ta}-\mathbb{E}_{\theta}[\mathbf{x}^{ta}|\mathbf{x}^{co}].$
Consequently, the target data $\mathbf{x}^{ta}$ for diffusion models in Eqs. (8), (2), and (3) is replaced by $\mathbf{r}^{ta}$. The residual distribution of urban spatiotemporal data is not independently and identically distributed (i.i.d.) nor does it follow a fixed distribution, such as $\mathcal{N}(0,\mathbf{\sigma})$. Instead, it often exhibits complex spatiotemporal dependence and heterogeneity. So we consider both temporal residual learning and spatial residual learning.
@@
Accordingly, the Eq (13) in the forward process is rewritten as:
@@
And in the denoising process, we sample $\mathbf{r}_{N}^{ta}$ from $\mathcal{N}({\mathbf{Q}},I)$, and denoise it use Eq (2), the computation of $\mu_{\theta}(\mathbf{r}_{n}^{ta},n|\mathbf{x}_{0}^{co})$ in Eq (2) is modified as:
@@ 5. Performance Evaluations
    DYffusion can only be applied to grid-format data.
@@ 5.1. Experimental Settings
To evaluate the performance, we employed two deterministic metrics, MAE and RMSE, along with three probabilistic metrics: CRPS, QICE and IS. For QICE, we set the number of QIs, denoted as $M_{\text{QIs}}$, to 10. For IS, we choose a confidence level of 90%, that is, $\alpha=0.1$.
@@ 5.3. Long-Term Prediction
Results of Deterministic Metrics.
@@ 5.5. Model Efficiency
We evaluated the training and time of our model and all baseline models on the CellularSH dataset, as well as the time required to perform 50 sampling iterations on the test set. Notably, the pretraining time of the mean prediction module in our model is also included in the evaluation. The experimental results are summarized in Table 5. As shown in Table 5, our model demonstrates a significant advantage over existing probabilistic models in terms of training and inference efficiency. This efficiency is particularly valuable for real-world deployment scenarios. It is worth noting that while the CSDI model benefits from the Transformer architecture’s capability to capture spatiotemporal dependencies, its training and sampling time is considerably higher than that of other models, posing challenges for time-sensitive tasks.
@@ 5.6. Analysis of Distribution Alignment
For regions with complex distributions in the CellularSH dataset, we validate the model’s ability to capture multi-modal characteristics by comparing the predicted and true distributions using Kernel Density Estimation (KDE). As shown in Figure 4, the real distribution exhibits significant spatiotemporal multi-modality: Figure 4(a) shows three peaks, which likely represent values at different time points and values under different states at the same time points. CoST successfully captures the three peaks, while CSDI only fits two, demonstrating superior capability in modeling multi-modal distributions. In Figure 4(b), the difference between the two peaks reflects strong temporal dynamics in this region. Although both models predict two peaks, the peak spacing of CoST is more consistent with the true distribution. These advantages stem from the model’s design: the diffusion model focuses on residuals, effectively capturing multi-modal distribution features, while the deterministic model captures long-term trends and enhances the model’s capacity to model primary dynamic patterns.